Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 141,000 kilometers during a year, the average operating cost is 12.1 cents per kilometer. If a truck is driven only 94,000 kilometers during a year, the average operating cost increases to 14.6 cents per kilometer.Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is drHoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is drHoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is dr
Answer #1
1
Kilometers Total cost
High level 141000 17061 =141000*0.121
Low level 94000 13724 =94000*0.146
Difference 47000 3337
Variable cost 0.071 per kilometer =3337/47000
Fixed cost 7050 per year =17061-(141000*0.071)
2
Y = 7050 + 0.071 X
3
Total annual cost 15357 =7050+(117000*0.071)